BexleyBrooks:

Solve the inequality and graph the solution on the line provided. 3x−7≤−22

BexleyBrooks:

* I don't have a line provided * but what would it be on a line?

Luigi0210:

The first step would be to actually solve the inequality, so you are given: \(\large 3x-7 \le -22 \) Add 7 to both sides \(\large 3x-7\color{red}{(+7)} \le -22\color{red}{(+7) }\) \(\large 3x \le -15 \) Now you can just divide by 3 to get x by itself \(\large 3x\color{red}{/3} \le -15 \color{red}{/3} \) \(\large x \le -5 \) Now you can plot this on a number line.